# Rankify: A Comprehensive Python Toolkit for Retrieval, Re-Ranking, and Retrieval-Augmented Generation

### v2

@@ 3.1. Prebuilt Retrieval Corpora and Indexes
Rankify provides two large-scale retrieval corpora: Wikipedia (Karpukhin et al., 2020) and MS MARCO (Nguyen et al., 2017), enabling researchers to conduct retrieval experiments on well-established benchmarks. To support efficient retrieval, Rankify includes pre-builtt index files for each retriever, allowing users to directly query the corpora without requiring costly indexing. Each corpus has been indexed for multiple retrieval models, including BM25, DPR, ANCE, BGE, Contriever, and ColBERT. By providing precomputed indexes, Rankify simplifies large-scale retrieval research while maintaining consistency across different retrieval pipelines. Users can view all available datasets and retrievers using the following command (see Listing 1):
@@ 3.2. Datasets
With Rankify, we aim to provide a standardized and efficient way to handle datasets for retrieval, re-ranking, and retrieval-augmented generation (RAG). Given the complexity of retrieval workflows, Rankify integrates pre-retrieved datasets with structured annotations, allowing users to experiment with different retrieval and re-ranking methods using just a few lines of code. Our dataset design unifies multiple annotation styles and storage formats, enabling a seamless interface for loading, processing, and benchmarking retrieval-based tasks.
The datasets included in Rankify fall into multiple categories, as illustrated in Figure 3, which showcases their distribution based on task type and dataset size. Each dataset in Rankify follows a standardized schema that consists of three main components: (1) Question: The input query for which relevant information needs to be retrieved. (2) Answers: The expected correct responses associated with the question. (3) Contexts (Or Retrieved Documents): A ranked list of candidate documents retrieved by a specific retriever. The Dataset class in Rankify encapsulates these attributes and provides easy access to retrieval, evaluation, and benchmarking. The dataset schema is depicted in Figure 4, illustrating how questions, answers, and contexts interact within Rankify.
To facilitate benchmarking and ease of use, Rankify provides pre-retrieved datasets where each query is associated with 1,000 retrieved documents from different retrievers. These datasets allow users to test different re-ranking and retrieval models without the need for time-consuming indexing.
As mentioned above, to ensure standardized retrieval benchmarking, Rankify provides pre-retrieved datasets, where each query is associated with 1,000 retrieved documents from Wikipedia. These datasets span multiple retrieval tasks, including question answering, multi-hop reasoning, temporal retrieval, entity linking, fact verification, dialogue generation, and summarization. To enable fair comparisons across different retrieval methods, each dataset has been processed using multiple retrievers, including BM25, DPR, ANCE, BGE, Contriever, and ColBERT. By applying the same retrieval pipeline across all the datasets, Rankify eliminates the need for computationally expensive indexing while maintaining consistency in evaluation.
@@
Listing 2: Loading a dataset in Rankify and inspecting its structure.
@@
5# Inspect the first document
6print(documents[0])
@@
2# Define a new dataset with a single question and answer
3question = Question("What is the capital of France?")
4answer = Answer(["Paris"])
5context = [Context(score=0.9, has_answer=True, id=1, title="France", text="The capital is Paris.")]
6# Create the document and dataset
7document = Document(question, answer, context)
8custom_dataset = Dataset(retriever="custom", dataset_name="MyDataset")
9custom_dataset.documents = [document]
10# Save the dataset
11custom_dataset.save_dataset("my_dataset.json")
@@
1from rankify.dataset.dataset import Dataset
2# Load a dataset that includes retrieved documents
3retrieval_dataset = Dataset.load_dataset("path/to/retrieval_dataset.json")
4# Load a QA dataset with only questions and answers
5qa_dataset = Dataset.load_dataset_qa("path/to/qa_dataset.jsonl")
@@ 3.3. Retriever Models
Rankify supports a diverse set of retriever models to facilitate document retrieval for various information retrieval tasks. The current version includes both sparse and dense retrievers. Sparse retrievers, such as BM25 (Robertson et al., 1995), rely on exact term matching, while dense retrievers, such as DPR (Yates et al., 2021), ANCE (Xiong et al., 2021a), and BGE (Chen et al., 2023), and ColBERT (Khattab and Zaharia, 2020).
The retrievers implemented in Rankify are accessible via a unified interface, where users can specify the retrieval method and parameters. The available retriever models include BM25, DPR, ANCE, Contriever (Izacard et al., 2021), BGB, and ColBERT. Rankify allows users to retrieve documents efficiently using different retrieval models. The retrievers can be applied to any dataset supported by Rankify, making it flexible for experimentation and benchmarking. Users can easily initialize a retriever and apply it to a list of query-document pairs, as demonstrated in Listing 5.
Listing 5: Using a retriever model in Rankify to retrieve documents.
@@ 8 Document(question=Question("Who wrote Hamlet?"),
In Listing 5, users can specify the retrieval corpus, select the desired retriever model, and define the number of documents to retrieve. This flexibility allows researchers to experiment with different retrieval settings, optimizing their retrieval strategy based on the task requirements.
@@ 3.4. Re-Ranking Models
In information retrieval, two-stage pipelines are widely used to maximize retrieval performance. The first stage involves retrieving a set of candidate documents using a computationally efficient retriever, followed by a second stage where these documents are re-ranked using a stronger, typically neural network-based, model. This re-ranking step enhances retrieval quality by considering deeper semantic relationships between the query and documents. While effective, re-ranking models vary significantly in their architectures, trade-offs, and implementations, making it challenging for users to select the most suitable method for their specific needs.
To address this, Rankify provides a unified interface for re-ranking, enabling users to effortlessly switch between different models with minimal modifications. The framework supports a diverse set of 24 primary re-ranking models with 41 sub-methods, spanning different re-ranking strategies such as pointwise, pairwise, and listwise approaches. The implemented models include MonoBERT (Nogueira and Cho, 2019), MonoT5 (Nogueira et al., 2020), RankT5 (Zhuang et al., 2023), ListT5 (Yoon et al., 2024), ColBERT (Santhanam et al., 2021), RankGPT (Sun et al., 2023b), and various transformer-based re-rankers.
@@ 3.5. Retrieval-Augmented Generation (RAG) Models
Rankify allows users to leverage large-scale language models such as LLaMA (Touvron et al., 2023), GPT-4 (Achiam et al., 2023), and T5-based models (Raffel et al., 2020) for retrieval-augmented generation. By supporting both encoder-decoder architectures (FiD (Izacard and Grave, 2020)) and decoder-only models (e.g., GPT, LLaMA), Rankify provides flexibility for optimizing generation quality based on task-specific requirements.
@@ 3.6. Evaluation Metrics
Rankify provides evaluation metrics for retrieval, re-ranking, and retrieval-augmented generation (RAG). For retrieval and re-ranking, Top-k accuracy measures whether relevant documents appear within the top-k results, allowing performance comparison before and after re-ranking. Listing 8 demonstrates how to compute Top-k accuracy.
@@ 4.1. Experiment Setup
Rankify enables researchers to benchmark retrieval, re-ranking, and RAG methods, evaluate their own approaches, and explore optimizations within these tasks. To demonstrate its capabilities, we conducted multiple experiments to provide reproducible benchmarks and performance insights. All main experiments were conducted on 8 NVIDIA A100 GPUs.
Experiments were performed on a diverse set of datasets, covering open-domain and multi-hop QA, fact verification, and temporal retrieval tasks. We evaluated retrieval performance on datasets such as Natural Questions (NQ), TriviaQA, HotpotQA, 2WikiMultiHopQA, and ArchivalQA, while re-ranking performance was analyzed on MSMARCO, WebQuestions, and PopQA. For retrieval evaluation, we measured Top-k accuracy at k=1, 5, 10, 20, 50, 100. Exact match (EM), Precision, Recall, Contains, and F1 were used as the primary evaluation metric for QA tasks.
@@
2WikiMultiHopQA (Ho et al., 2020) dev 17.17 31.72 40.14 47.73 55.88 61.17 ArchivalQA (Wang et al., 2022a) dev 18.18 33.29 39.72 45.96 53.68 58.74
train 8.85 19.78 28.15 38.16 51.16 57.70 test 17.68 32.55 39.26 45.36 52.76 58.08
ChroniclingAmericaQA (Piryani et al., 2024) test 4.18 11.23 15.83 20.87 27.49 32.48 AmbigQA (Min et al., 2020; Kwiatkowski et al., 2019) train 27.30 50.67 59.63 67.63 75.68 79.87
dev 4.26 11.80 16.42 21.50 28.19 33.15 dev 30.22 54.25 65.58 72.98 80.47 84.92
ARC (Clark et al., 2018) test 4.96 19.84 33.57 51.94 79.09 91.80 TriviaQA (Joshi et al., 2017) dev 48.70 67.57 72.84 77.38 81.19 83.59
dev 4.83 19.45 30.84 52.13 79.06 91.83 test 48.22 67.44 72.79 77.28 81.42 83.87
train 4.45 18.16 31.39 51.39 78.19 91.36 train 47.87 67.22 72.90 77.15 81.30 83.64
BoolQ (Clark et al., 2019) train 1.71 4.33 7.23 12.03 24.56 37.76 CommenseQA (Talmor et al., 2019) train 2.85 8.74 13.36 19.28 28.87 36.54
dev 1.22 3.94 6.61 11.68 23.67 37.49 dev 2.13 6.96 11.88 17.77 27.11 33.82
CuratedTREC (Lee et al., 2020) train 16.98 29.77 33.95 36.74 40.00 40.93 AY2 (Hoffart et al., 2011) train 23.81 41.60 48.84 55.88 63.78 69.62
test 18.14 31.63 34.65 38.84 41.86 43.02 dev 25.02 41.14 48.10 55.39 64.19 70.55
FEVER (Thorne et al., 2018; Petroni et al., 2021) dev 0.05 0.41 0.90 1.70 3.78 6.50 HellaSwag (Zellers et al., 2019) train 3.97 16.60 28.86 46.02 71.68 85.20
train 0.15 0.74 1.42 2.52 5.85 10.49 dev 3.77 15.00 26.31 42.89 69.00 83.51
SQuAD (Rajpurkar et al., 2016) dev 36.94 57.61 64.78 71.31 77.89 81.96 MMLU (Hendrycks et al., 2021b, a) train 3.65 14.14 23.98 38.68 62.22 78.04
test 36.67 57.37 64.50 71.09 77.74 81.84 dev 6.53 21.75 33.64 50.62 75.77 88.18
train 31.84 50.24 57.15 63.33 70.32 74.61 test 6.53 22.01 34.67 51.22 74.35 86.40
QuaRTz  (Tafjord et al., 2019) test 5.87 23.60 38.90 58.67 80.74 90.43 NarrativeQA (s Koˇciský et al., 2018) dev 15.46 23.92 27.10 30.45 35.08 38.89
dev 7.81 26.82 40.62 63.54 84.38 92.19 test 16.80 25.24 28.46 31.50 35.74 38.65
train 7.81 26.82 40.62 63.54 84.38 92.19 train 16.80 25.24 28.46 31.50 35.74 38.65
NQ (Kwiatkowski et al., 2019) dev 23.62 45.60 55.41 64.19 72.51 77.72 OpenBookQA (Mihaylov et al., 2018) dev 3.40 15.20 31.60 50.60 79.00 90.00
test 23.46 45.62 56.32 64.88 74.57 79.72 train 3.35 16.24 28.99 47.85 75.01 88.56
train 22.89 44.97 54.32 62.73 71.75 77.03 test 4.00 16.80 32.20 51.20 74.00 88.80
WebQuestions (Berant et al., 2013) test 19.54 42.86 53.44 63.24 72.34 76.43 Musique (Trivedi et al., 2022) dev 7.36 15.47 21.39 25.98 34.59 41.29
train 21.41 44.71 55.35 64.21 72.71 78.45 train 6.23 13.70 18.54 24.39 33.77 40.81
T-REx (ElSahar et al., 2018; Petroni et al., 2021) dev 45.48 64.34 70.60 76.04 80.42 82.98 TruthfulQA (Lin et al., 2022b) dev 2.33 7.59 12.12 20.69 38.19 56.55
SIQA (Sap et al., 2019) train 0.34 1.52 2.78 4.85 8.92 13.43 HotpotQA (Yang et al., 2018) dev 28.13 45.02 51.92 57.79 65.44 70.67
dev 0.61 2.05 3.58 6.55 11.31 16.58 train 34.89 51.71 58.13 64.09 71.11 75.49
WOW (Dinan et al., 2019; Petroni et al., 2021) dev 0.20 0.33 0.36 0.39 0.46 0.46 ZSRC (Levy et al., 2017) dev 52.26 70.86 76.29 81.44 85.39 87.73
train 0.26 0.38 0.41 0.45 0.49 0.52 train 50.53 67.23 72.06 76.10 80.35 82.99
Table 2. Retrieval performance of BM25 across several selected datasets. The table reports Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy for each dataset split. Datasets vary in complexity, including open-domain QA (e.g., TriviaQA, NQ), multi-hop reasoning (e.g., 2WikiMultiHopQA, HotpotQA), fact verification (e.g., FEVER, TruthfulQA), and temporal retrieval (e.g., ArchivalQA, ChroniclingQA).
4.2. Retrieval Results
The retrieval performance of BM25 across multiple datasets is shown in Table 2. The results report Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy for each dataset split. Performance varies significantly based on dataset characteristics, with open-domain QA datasets such as TriviaQA and Natural Questions achieving high Top-k accuracy, while more complex multi-hop datasets like 2WikiMultiHopQA and HotpotQA show lower retrieval rates, reflecting the challenges in retrieving supporting evidence across multiple documents. Temporal datasets like ArchivalQA and ChroniclingQA exhibit moderate retrieval performance, likely due to the temporal nature of their content.
For datasets requiring precise fact verification, such as FEVER and TruthfulQA, BM25 struggles, achieving low Top-1 accuracy, indicating the limitations of keyword-based retrieval in distinguishing fine-grained factual nuances. In contrast, reasoning-based datasets like StrategyQA and CommonsenseQA show relatively low retrieval accuracy, highlighting the need for more advanced retrievers capable of capturing implicit reasoning dependencies.
Table 3 presents the retrieval performance of various retrievers across three benchmark datasets: NQ, WebQ, and TriviaQA. The results report Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy, providing insights into how different retrieval models rank relevant documents. Among the tested retrievers, dense retrieval methods such as DPR, MSS-DPR, and ColBERT consistently outperform retrievers like MSS and Contriever. MSS-DPR achieves the highest Top-1 accuracy, reaching 50.16% on NQ, 44.24% on WebQ, and 61.63% on TriviaQA, demonstrating the effectiveness of fine-tuned dense retrieval. Standard DPR also performs well, particularly on NQ and TriviaQA, where it reaches 48.67% and 57.47% Top-1 accuracy, respectively. ColBERT, which employs late-interaction ranking, shows competitive performance, particularly in higher recall settings (Top-50 and Top-100).
@@ 4.2. Retrieval Results
MSS (Sachan et al., 2021) NQ 19.27 41.24 51.27 59.97 69.55 75.56
WebQ 11.66 29.03 39.12 49.21 61.12 68.35
TriviaQA 30.76 52.64 60.52 67.17 74.57 79.11
MSS-DPR (Sachan et al., 2021) NQ 50.16 71.88 77.47 81.44 85.98 88.14
WebQ 44.24 65.00 71.65 76.91 81.84 84.54
TriviaQA 61.63 75.21 79.14 81.85 84.92 86.58
Contriever (Izacard et al., 2021) NQ 22.16 47.28 58.72 67.86 76.01 80.55
WebQ 19.98 43.45 56.39 65.69 74.85 80.11
TriviaQA 34.16 59.48 68.00 73.91 79.83 82.93
DPR (Yates et al., 2021) NQ 48.67 68.78 74.54 79.19 83.71 85.70
WebQ 44.83 65.00 70.62 74.60 78.69 81.64
@@
Table 3. Comparison of different retrieval models on Natural Questions (NQ), WebQuestions (WebQ), and TriviaQA. The table reports Top-1 to Top-100 accuracy for each retriever.
Reranking/ Model NQ-Test WebQ
Top-1 Top-5 Top-10 Top-20 Top-50 Top-1 Top-5 Top-10 Top-20 Top-50
BM25 - 22.10 43.76 54.45 62.93 72.74 18.80 41.83 52.16 62.40 71.70
UPR (Sachan et al., 2022) T5-small 23.60 49.44 59.97 67.70 75.15 18.55 43.65 55.56 64.91 72.68
T5-base 26.81 52.93 63.32 69.77 76.12 20.66 48.52 58.56 66.48 72.68
T5-large 29.75 55.84 65.67 72.02 76.48 24.21 51.77 60.38 67.27 73.32
T0-3B 35.42 60.44 67.56 72.68 76.75 32.48 56.39 64.17 69.09 73.67
gpt-neo-2.7B 28.75 55.01 64.81 71.13 76.56 24.75 48.91 59.64 66.83 72.63
gpt2 25.95 50.08 60.47 69.05 75.87 20.47 45.32 56.49 65.69 72.78
gpt2-medium 26.75 52.99 63.04 69.80 75.95 22.39 49.01 59.54 66.92 72.44
gpt2-large 26.59 52.35 62.68 70.13 75.95 24.06 49.85 59.84 67.07 72.78
gpt2-xl 27.28 53.85 63.24 70.11 75.84 23.57 50.78 60.48 66.87 72.73
RankGPT (Sun et al., 2023a) llamav3.1-8b 41.55 60.30 66.17 70.60 75.42 38.77 58.36 62.69 67.71 73.12
FlashRank (Damodaran, 2024) TinyBERT-L-2-v2 31.49 53.90 61.57 68.33 74.95 28.54 51.87 60.62 67.42 73.17
MiniLM-L-12-v2 41.02 62.07 67.50 72.49 76.26 37.05 58.07 64.96 70.02 73.96
MultiBERT-L-12 11.99 31.74 43.54 55.01 69.63 12.54 34.30 45.91 55.90 67.91
ce-esci-MiniLM-L12-v2 34.70 57.47 64.81 70.60 76.17 31.84 55.01 62.54 68.60 73.47
RankT5 (Zhuang et al., 2023) base 43.04 62.99 68.47 72.99 76.28 36.95 58.75 64.27 70.07 74.45
large 45.54 64.98 70.02 73.85 76.81 38.77 60.28 66.48 70.96 74.31
3b 47.17 66.45 70.85 74.18 76.89 40.40 60.33 66.58 71.16 74.45
Inranker (Laitz et al., 2024) small 15.90 40.11 46.84 56.23 69.83 14.46 36.36 46.25 57.48 69.98
base 15.90 40.69 48.11 56.62 69.66 14.46 37.45 46.80 56.54 69.68
3b 15.90 40.22 48.06 57.97 69.00 14.46 37.30 46.11 57.48 69.34
LLM2Vec (BehnamGhader et al., 2024) Meta-Llama-31-8B 24.32 49.22 59.55 68.08 75.26 26.72 50.39 60.48 68.40 73.47
MonoBert (Nogueira et al., 2019) large 39.05 61.71 67.89 72.63 76.56 34.99 57.33 64.56 70.12 73.96
Twolar (Baldelli et al., 2024) twolar-xl 46.84 65.29 70.22 73.62 76.86 41.68 60.82 67.07 71.55 74.40
Echorank (Rashid et al., 2024) flan-t5-large 36.73 55.09 59.11 62.38 62.38 31.74 53.69 58.75 61.51 61.51
flan-t5-xl 41.68 55.15 59.05 62.38 62.38 36.22 52.55 57.18 61.51 61.51
Incontext Reranker (Chen et al., 2024) llamav3.1-8b 15.15 38.50 57.11 69.66 76.48 18.89 41.83 52.16 62.40 71.70
Lit5 (Tamber et al., 2023) LiT5-Distill-base 40.05 59.80 65.95 70.80 75.73 36.76 56.79 63.48 67.81 73.12
LiT5-Distill-large 44.40 62.57 67.59 71.74 76.01 39.66 58.02 64.56 68.94 73.67
LiT5-Distill-xl 47.81 64.07 68.55 72.07 76.26 42.37 59.44 65.55 69.38 73.62
LiT5-Distill-base-v2 42.57 61.49 66.73 70.83 75.56 39.61 57.97 64.22 68.55 73.32
LiT5-Distill-large-v2 46.53 63.10 67.83 71.91 75.87 41.97 59.49 65.64 68.65 72.98
LiT5-Distill-xl-v2 47.92 64.18 69.03 72.40 76.17 41.53 60.13 65.69 69.04 73.27
Sentence Transformer Reranker all-MiniLM-L6-v2 (Wang et al., 2020) 33.35 57.34 65.37 70.72 76.01 30.95 53.44 62.10 68.01 73.52
GTR-base (Ni et al., 2021) 39.41 59.08 65.95 71.32 76.03 36.56 58.21 64.32 69.14 73.62
GTR-large 40.63 62.07 68.25 72.32 76.73 38.97 58.51 65.30 70.02 73.57
GTR-xl 41.55 61.82 67.78 72.65 76.81 38.92 59.69 66.04 70.57 74.01
GTR-xxl 42.93 63.32 68.55 73.01 77.00 39.41 60.03 65.89 70.47 74.01
T5-base (Raffel et al., 2020) 31.19 55.84 63.60 70.22 76.06 29.77 54.92 62.84 68.70 73.52
T5-xxl 38.89 61.46 67.78 72.49 76.64 35.82 58.26 65.206 70.32 74.01
T5-large 30.80 55.59 63.35 70.02 76.37 30.51 55.06 61.71 68.35 73.37

### v3

@@ 3.1. Prebuilt Retrieval Corpora and Indexes
Rankify provides two large-scale retrieval corpora: Wikipedia (Karpukhin et al., 2020) and MS MARCO (Nguyen et al., 2017), enabling researchers to conduct retrieval experiments on well-established benchmarks. To support efficient retrieval, Rankify includes pre-built index files for each retriever, allowing users to directly query the corpora without requiring costly indexing. Each corpus has been indexed for multiple retrieval models, including BM25, DPR, ANCE, BGE, Contriever, and ColBERT. By providing precomputed indexes, Rankify simplifies large-scale retrieval research while maintaining consistency across different retrieval pipelines. Users can view all available datasets using the following command (see Listing 1):
@@ 3.2. Datasets
Rankify provides a standardized and efficient way to handle datasets for retrieval, re-ranking, and retrieval-augmented generation (RAG). To simplify retrieval workflows, Rankify integrates pre-retrieved datasets with structured annotations, enabling users to experiment with various retrieval and re-ranking methods with minimal code. The datasets in Rankify are categorized based on task type and size, as shown in Table 2. Each dataset follows a standardized schema with three main components: the Question, which represents the query requiring relevant information; the Answers, which are the expected correct responses to the query; and the Contexts (or Retrieved Documents), which provide a ranked list of candidate documents retrieved by a specific retriever.
To support benchmarking and reproducibility, Rankify provides pre-retrieved datasets, each with 1,000 top-ranked documents from Wikipedia [^8] Note: We are currently processing different retrievers based on MS MARCO to generate and store 1,000 top-ranked documents per query for each dataset., processed by various retrievers such as BM25, DPR, ANCE, BGE, Contriever, and ColBERT. These datasets cover diverse Re-Ranking and RAG tasks.
@@
Listing 2: Loading a dataset in Rankify and inspecting its structure.
@@
2question = Question("What is the capital of France?")
3answer = Answer(["Paris"])
4context = [Context(score=0.9, has_answer=True, id=1, title="France", text="The capital is Paris.")]
5document = Document(question, answer, context)
6custom_dataset = Dataset(retriever="custom", dataset_name="MyDataset")
7custom_dataset.documents = [document]
8custom_dataset.save_dataset("my_dataset.json")
@@
1from rankify.dataset.dataset import Dataset
2retrieval_dataset = Dataset.load_dataset("path/to/retrieval_dataset.json")
3qa_dataset = Dataset.load_dataset_qa("path/to/qa_dataset.jsonl")
@@ 3.3. Retriever Models
Rankify supports a diverse set of retriever models to facilitate document retrieval for various information retrieval tasks. The current version includes both sparse retrievers, such as BM25 (Robertson et al., 1995), and dense retrievers, such as DPR (Yates et al., 2021), ANCE (Xiong et al., 2021), and BGE (Chen et al., 2023), Contriever (Izacard et al., 2021a) and ColBERT (Khattab and Zaharia, 2020). The retrievers implemented in Rankify are accessible via a unified interface, where users can specify the retrieval method and parameters.
Rankify allows users to retrieve documents efficiently using different retrieval models. The retrievers can be applied to any dataset supported by Rankify, making it flexible for experimentation and benchmarking. Users can easily initialize a retriever and apply it to a list of query-document pairs, as demonstrated in Listing 5.
@@ 8 Document(question=Question("Who wrote Hamlet?"),
Listing 5: Using a retriever model in Rankify to retrieve documents.
In Listing 5, users can specify the retrieval corpus by setting index_type to msmarco or wiki, select the desired retriever model, and define the number of documents to retrieve. This flexibility allows researchers to experiment with different retrieval settings, optimizing their retrieval strategy based on the task requirements.
@@ 3.4. Re-Ranking Models
In information retrieval, two-stage pipelines are widely used to maximize retrieval performance. The first stage involves retrieving a set of candidate documents using a computationally efficient retriever, followed by a second stage where these documents are re-ranked using a stronger, typically, neural network-based model. This re-ranking step enhances retrieval quality by considering deeper semantic relationships between the query and documents. While effective, re-ranking models vary significantly in their architectures, trade-offs, and implementations, making it challenging for users to select the most suitable method for their specific needs.
To address this, Rankify provides a unified interface for re-ranking, enabling users to effortlessly switch between different models with minimal modifications. The framework supports a diverse set of 24 primary re-ranking models with 41 sub-methods, spanning different re-ranking strategies such as pointwise, pairwise, and listwise approaches. The implemented models include MonoBERT (Nogueira and Cho, 2019), MonoT5 (Nogueira et al., 2020), RankT5 (Zhuang et al., 2023), ListT5 (Yoon et al., 2024), ColBERT (Santhanam et al., 2021), RankGPT (Sun et al., 2023), and various transformer-based re-rankers.
@@ 3.5. Retrieval-Augmented Generation (RAG) Models
Rankify allows users to leverage large-scale language models such as LLaMA (Touvron et al., 2023), GPT-4 (Achiam et al., 2023), and T5-based models (Raffel et al., 2020) for retrieval-augmented generation. By supporting both encoder-decoder architectures (FiD (Izacard and Grave, 2020)) and decoder-only models (e.g., GPT, LLaMA), the framework provides flexibility for optimizing generation quality based on task-specific requirements.
@@ 3.6. Evaluation Metrics
Table 3. Retrieval performance of BM25 across several selected datasets. The table reports Top-1, Top-5, Top-10, Top-20, Top-50, and Top-100 retrieval accuracy for each dataset split. Datasets vary in complexity, including open-domain QA (e.g., TriviaQA, NQ), multi-hop reasoning (e.g., 2WikiMultiHopQA, HotpotQA), fact verification (e.g., TruthfulQA), and temporal retrieval (e.g., ArchivalQA, ChroniclingAmericaQA).
2WikiMultiHopQA (Ho et al., 2020) Train 8.85 19.78 28.15 38.16 51.16 57.70 ArchivalQA (Wang et al., 2022a) Val 18.18 33.29 39.72 45.96 53.68 58.74
Val 17.17 31.72 40.14 47.73 55.88 61.17 Test 17.68 32.55 39.26 45.36 52.76 58.08
ChroniclingAmericaQA (Piryani et al., 2024) Val 4.26 11.80 16.42 21.50 28.19 33.15 AmbigQA (Min et al., 2020; Kwiatkowski et al., 2019) Train 27.30 50.67 59.63 67.63 75.68 79.87
Test 4.18 11.23 15.83 20.87 27.49 32.48 Val 30.22 54.25 65.58 72.98 80.47 84.92
ARC (Clark et al., 2018) Train 4.45 18.16 31.39 51.39 78.19 91.36 TriviaQA (Joshi et al., 2017) Train 47.87 67.22 72.90 77.15 81.30 83.64
Val 4.83 19.45 30.84 52.13 79.06 91.83 Val 48.70 67.57 72.84 77.38 81.19 83.59
Test 4.96 19.84 33.57 51.94 79.09 91.80 Test 48.22 67.44 72.79 77.28 81.42 83.87
SQuAD (Rajpurkar et al., 2016) Train 31.84 50.24 57.15 63.33 70.32 74.61 MMLU (Hendrycks et al., 2021b, a) Train 3.65 14.14 23.98 38.68 62.22 78.04
Val 36.94 57.61 64.78 71.31 77.89 81.96 Val 6.53 21.75 33.64 50.62 75.77 88.18
Test 36.67 57.37 64.50 71.09 77.74 81.84 Test 6.53 22.01 34.67 51.22 74.35 86.40
QuaRTz (Tafjord et al., 2019) Train 7.81 26.82 40.62 63.54 84.38 92.19 NarrativeQA (s Koˇciský et al., 2018) Train 16.80 25.24 28.46 31.50 35.74 38.65
Val 7.81 26.82 40.62 63.54 84.38 92.19 Val 15.46 23.92 27.10 30.45 35.08 38.89
Test 5.87 23.60 38.90 58.67 80.74 90.43 Test 16.80 25.24 28.46 31.50 35.74 38.65
NQ (Kwiatkowski et al., 2019) Train 22.89 44.97 54.32 62.73 71.75 77.03 OpenBookQA (Mihaylov et al., 2018) Train 3.35 16.24 28.99 47.85 75.01 88.56
Val 23.62 45.60 55.41 64.19 72.51 77.72 Val 3.40 15.20 31.60 50.60 79.00 90.00
Test 23.46 45.62 56.32 64.88 74.57 79.72 Test 4.00 16.80 32.20 51.20 74.00 88.80
WebQuestions (Berant et al., 2013) Train 21.41 44.71 55.35 64.21 72.71 78.45 Musique (Trivedi et al., 2022) Train 6.23 13.70 18.54 24.39 33.77 40.81
Test 19.54 42.86 53.44 63.25 72.34 76.43 Val 7.36 15.47 21.39 25.98 34.59 41.29
T-REx (ElSahar et al., 2018; Petroni et al., 2021) Val 45.48 64.34 70.60 76.04 80.42 82.98 TruthfulQA (Lin et al., 2022) Val 2.33 7.59 12.12 20.69 38.19 56.55
SIQA (Sap et al., 2019) Train 0.34 1.52 2.78 4.85 8.92 13.43 HotpotQA (Yang et al., 2018) Train 34.89 51.71 58.13 64.09 71.11 75.49
Val 0.61 2.05 3.58 6.55 11.31 16.58 Val 28.13 45.02 51.92 57.79 65.44 70.67
WOW (Dinan et al., 2019; Petroni et al., 2021) Train 0.26 0.38 0.41 0.45 0.49 0.52 ZSRC (Levy et al., 2017) Train 50.53 67.23 72.06 76.10 80.35 82.99
Val 0.20 0.33 0.36 0.39 0.46 0.46 Val 52.26 70.86 76.29 81.44 85.39 87.73
Rankify provides evaluation metrics for retrieval, re-ranking, and retrieval-augmented generation (RAG). For retrieval and re-ranking, Top-k accuracy measures whether documents containing the correct answer appear within the top-k results. This definition of relevance, while task-specific, aligns with the goal of downstream applications like question answering. Listing 8 demonstrates how to compute Top-k accuracy.
@@ 4.1. Experiment Setup
Rankify enables researchers to benchmark retrieval, re-ranking, and RAG methods, evaluate their own approaches, and explore optimizations within these tasks. To demonstrate its capabilities, we conducted multiple experiments to provide reproducible benchmarks and performance insights. All main experiments were conducted on 2x NVIDIA A100 GPUs.
Experiments were performed on a diverse set of datasets, covering open-domain and multi-hop QA, fact verification, and temporal retrieval tasks. All the experiments in our study use preprocessed English Wikipedia dump from December 2018 as released by (Yates et al., 2021) as evidence passages. Each Wikipedia article is split into non-overlapping 100-word passages, with over 21 million passages in total.
We evaluated retrieval performance on datasets such as Natural Questions (NQ), TriviaQA, HotpotQA, 2WikiMultiHopQA, and ArchivalQA, while re-ranking performance was analyzed on MSMARCO, WebQuestions, and PopQA. For retrieval evaluation, we measured Top-k accuracy at k=1, 5, 10, 20, 50, 100. Exact match (EM), Precision, Recall, Contains, and F1 were used as the primary evaluation metrics for QA tasks.
@@
Table 4. Comparison of different retrieval models on NQ, WebQ, and TriviaQA. The table reports Top-1 to Top-100 accuracy for each retriever.
@@
Contriever (Izacard et al., 2021a) NQ 38.81 65.65 73.91 79.56 84.88 88.01
WebQ 35.97 63.83 69.78 75.94 81.64 83.76
TriviaQA 49.85 71.39 76.68 80.26 83.85 85.72
DPR (Yates et al., 2021) NQ 44.57 67.76 74.52 79.50 84.40 86.81
WebQ 44.64 63.98 70.52 75.05 80.51 82.97
@@
Ance (Xiong et al., 2021) NQ 50.80 71.86 78.12 82.52 86.23 88.28
WebQ 46.51 66.24 71.80 76.62 81.79 84.25
TriviaQA 56.46 72.14 76.65 80.04 83.37 85.22
BGE (Xiao et al., 2023; Chen et al., 2023) NQ 48.03 72.22 78.50 82.66 86.90 89.45
WebQ 42.67 65.45 72.54 79.04 82.78 85.88
TriviaQA 57.81 75.39 79.83 82.92 85.53 86.94
Table 5. Top-20/Top-100 retrieval accuracy of different retrievers on three open-domain QA datasets (NQ, TriviaQA, and WebQ). Our Rankify implementation achieves results identical to Pyserini across all retrievers, as we use Pyserini as the indexing backend. This validates the correctness of our implementation. We also match the official GitHub results for ColBERT and Contriever on these datasets.
Retrievers NQ TriviaQA WebQ
DPR (Yates et al., 2021) 78.4 / 85.4 79.4 / 85.0 73.2 / 81.4
DPR_Pyserini (Lin et al., 2021) 79.5 / 86.8 79.7 / 85.1 75.1 / 82.9
DPR_Rankify 79.5 / 86.8 79.7 / 85.1 75.1 / 82.9
BM25 (Sachan et al., 2022) 63.0 / 78.2 76.4 / 83.1 62.3 / 75.5
BM25_Pyserini (Lin et al., 2021) 64.8 / 79.7 77.3 / 83.8 63.3 / 76.4
BM25_Rankify 64.8 / 79.7 77.3 / 83.8 63.3 / 76.4
Contriever (Izacard et al., 2021b) 79.6 / 88.0 80.4 / 85.7 75.9 / 83.7
Contriever_Rankify 79.6 / 88.0 80.3 / 85.7 75.9 / 83.7
ColBert (Santhanam et al., 2021) 82.1 / 88.5 82.2 / 86.4 76.6 / 84.5
ColBert_Rankify 82.1 / 88.5 82.2 / 86.4 76.6 / 84.5
ANCE (Xiong et al., 2021) 82.1 / 87.9 80.3 / 85.2 - / -
ANCE_Rankify 82.5 / 88.5 80.0 / 85.2 76.6 / 84.25